OpenStudy (anonymous):

6.5X+1.4y=59 What is the value of x when y=5? i dont know how to do this, help??

OpenStudy (anonymous):

you can look at this equation like this, 1.4 is being multiplied to the y right?

OpenStudy (anonymous):

you can see that?

OpenStudy (anonymous):

Yess

OpenStudy (anonymous):

so that y in that equation is just a place holder. you can plug in any number into that y and see what x is

OpenStudy (anonymous):

So would you multiply the x to 5?

OpenStudy (anonymous):

for example x + y = 3 what is x when y = 2 you would replace 2 for y ===> x + 2 = 3 now you can solve for x

OpenStudy (anonymous):

I'm so confused, I'm sorry, I'm not good in math lol. you're gonna have to break it down more, if possible.

OpenStudy (anonymous):

for this one you have 6.5x + 1.4y = 59 what is x when y = 5 so replace y with 5 6.5x + 1.4 * 5 = 59

OpenStudy (anonymous):

now you can solve for x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay, haha

OpenStudy (anonymous):

let me know what is confusing to you. cause the equation can look confusing because of the decimals.

OpenStudy (anonymous):

Like, I understand how to replace it now, but what am I supposed to do??

OpenStudy (anonymous):

okay great

OpenStudy (anonymous):

so can i ask you to show me what you get

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

like what does it look like when you replace that 5 with that y. the y should be gone and the 5 should be in its place.

OpenStudy (anonymous):

yes got that part

OpenStudy (anonymous):

can you show me how it looks by typing it in

OpenStudy (anonymous):

your going to be a champion at this after no worries. because these problems are the same everytime

OpenStudy (anonymous):

6.5x+5=59 And thanks :)

OpenStudy (anonymous):

yeah that looks good. but remember you are multiplying by that 1.4

OpenStudy (anonymous):

because the y was multiplied by the 1.4 the 5 should also be multiplied by 1.4

OpenStudy (anonymous):

so, 6.5x+7=59 ?

OpenStudy (anonymous):

and then do you subract the 7 from both sides, and divide by 6.5?

OpenStudy (anonymous):

yeah thats perfect. you seem to be good at solving for x. because waht you just said is exactly how you "SOLVE FOR X" is what its called lol just take a step back and look at what you did. you plugged 5 into y so it went from 6.5x + 1.4y = 59 to 6.5x + 1.4(5) = 59 and yeah the 1.4(5) = 7 so its exactly what you got 6.5x + 7 = 59

OpenStudy (anonymous):

YAYYY thank you! Um, could you also help me with something else?

OpenStudy (anonymous):

what you did is called "plugging in" you stuck the 5 into the y. and that means when y =5 x = when you solve for x like you said

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

but i need help with another problem

OpenStudy (anonymous):

yeah def

OpenStudy (anonymous):

okay, so i have -7(x-2)+5(-3x)-4x I know how to multiply them while in the parenthesis, but how do i solve for like terms?

OpenStudy (anonymous):

like i said these problems are the same everytime you plug in a number into x OR y what ever they ask. and then you can find out what the other one is

OpenStudy (anonymous):

okay yeah multiply them out and type what you get

OpenStudy (anonymous):

i screwed it up

OpenStudy (anonymous):

it was +5(3-X)

OpenStudy (anonymous):

i got -7x-14+15-5x-4x

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

now what do i do?

OpenStudy (anonymous):

okay you can think of the x's as money.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when its negative you owe a friend some money you borrowed and when its positive you owed money

OpenStudy (anonymous):

looks like you may be broke lol just because of the equation. so what you do is you all the like terms

OpenStudy (anonymous):

so you can subtract -7-5-4

OpenStudy (anonymous):

alright..

OpenStudy (anonymous):

it is -16. also bring the other ones together you have -14+15 = 1

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

so the final result is -16x + 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OHH got it (:

OpenStudy (anonymous):

it looks nice like this 1 - 16x

OpenStudy (anonymous):

right? lol

OpenStudy (anonymous):

but either way is correct

OpenStudy (anonymous):

thank you so so so muchh(:

OpenStudy (anonymous):

yeah i think you learned a lot

OpenStudy (anonymous):

you learned to plug in

OpenStudy (anonymous):

the second thing is called "accumulating terms" a term is something in an equation that is being added or subtracted

OpenStudy (anonymous):

so in your equation its 5 terms

OpenStudy (anonymous):

like 2x + 3 has 2

OpenStudy (anonymous):

also this is a term 2(x-4)

OpenStudy (anonymous):

so it could be 2(x-4) + 4(x-5) ^ ^ those are both terms so theres two

OpenStudy (anonymous):

okay i gtg peace

OpenStudy (anonymous):

can you give me a medal lol

OpenStudy (anonymous):

lol how

OpenStudy (anonymous):

next to my name on the right it says give medal

OpenStudy (anonymous):

i gave you one

OpenStudy (anonymous):

next to my name on the right it says give medal

OpenStudy (anonymous):

thanks(: okay

OpenStudy (anonymous):

you didnt give me on! lol no biggy. are you in highschool?

OpenStudy (anonymous):

one*